OpenStudy (anonymous):

Choose the value of the coefficient determinant (D) in the following system. 2x – 5y = –2 x – 3y = 0 –11 –1 1 11

OpenStudy (anonymous):

i was explaind to u before .. can u calculate of : 2(-3) - (-5)(1) = .... ?

OpenStudy (anonymous):

6-(-5)

OpenStudy (anonymous):

?

OpenStudy (anonymous):

its 11

OpenStudy (anonymous):

2(-3) - (-5)(1) = -6 - (-5) = -6 + 5 = -1

OpenStudy (anonymous):

so in this one x – 2y = –4 2x – 3y = –3 1-(-3)-2*2

OpenStudy (anonymous):

oh gosh am really bad sorry ! :(